The revenue from selling x units of a product is given by y = -0.0002x2 + 20x. How many units must be sold in order to have the greatest revenue? (Find the x-coordinate of the vertex of the parabola.)


Answer by stanbon(75887)

The revenue from selling x units of a product is given by y = -0.0002x2 + 20x. How many units must be sold in order to have the greatest revenue? (Find the x-coordinate of the vertex of the parabola.)
------
Max occurs when x = -b/2a
---
x = -20/(2*-0.0002) = 50,000
====================================
Cheers,
Stan H.